POLARIZATION QUANTUM BEAT SPECTROSCOPY OF HCF (A^{1}A′′): ^{19}F AND ^{1}H NUCLEAR HYPER-FINE STRUCTURE AND THE ZEEMAN EFFECT

\underline{J. XIN}, Department of Physics and Engineering Technologies, Bloomsburg University, Bloomsburg, PA 17815; H. FAN, I. IONESCU, C. ANNESLEY, AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233.

To investigate the ^{19}F and ^{1}H nuclear hyperfine structure and Zeeman effect in the simplest singlet carbene, HCF, we recorded polarization quantum beat spectra (QBS) of ^{r}R_{0}(J) lines of the pure bending transitions 2_{0}^{n} and combination bands 1_{0}^{1}2_{0}^{n} and 2_{0}^{n}3_{0}^{1} in the HCF A^{1}A′′-X^{1}A′ system. The spectra were measured under jet-cooled conditions using a pulsed discharge source, both at zero-field and under application of a weak magnetic field (< 30 G). Analysis yielded the nuclear spin-rotation constants (C_{aa}) and weak field Lande g_{aa} factors. Consistent with theoretical expectations, the majority of vibrational levels exhibit a linear correlation of C_{aa} and g_{aa}, from which the a hyperfine constants were estimated. The derived values are consistent with expectations based on the known hyperfine constants of related molecules.